OpenStudy (anonymous):

Find the second derivative y= x^2lnx

OpenStudy (anonymous):

Here is y' \[y'= 2 x \ln x + x^2 \frac 1 x= 2 x \ln x +x \] Try to do y''

OpenStudy (anonymous):

I use the product rule in my previous post.

OpenStudy (anonymous):

The second derivative of y=x^2lnx is 2lnx+3.

OpenStudy (anonymous):

Ive gotten to 2ln + (1/x+1)*2x i dont see how that simplifies to 3 could someone explain?

myininaya (myininaya):

\[y''=(2x \ln(x)+x)'=(2x \ln(x) )' +(x)'=2(x \ln(x))'+1=2(1 \ln(x)+x \frac{1}{x})+1\]

myininaya (myininaya):

I used the following rules: (f+g)'=f'+g' (fg)'=f'g+fg' (cf)'=cf' or we can even say (cfg)'=c(fg)'=c(f'g+fg')